OpenStudy (anonymous):

y varies directly as the square of x. When x = 2, y = 20. Find y when x = 4.

OpenStudy (anonymous):

y= x^2+16 now as x= 2 then y=2^2 +16 = 20 as x=4 then y= 4^2 +16 = 32

OpenStudy (anonymous):

well done!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

welcome